OpenStudy (anonymous):

Check all that apply. 3x2 + 7x + 8 = 6 A. x = 6 and x = -2 B. x = 6 and x = -1 C. x = -1/3 and x = -2 D. x = 3 and x = -2

OpenStudy (anonymous):

C

OpenStudy (anonymous):

how did you get that may i ask?

OpenStudy (anonymous):

Umm you can't ask

OpenStudy (anonymous):

:I fine then answer this one. please? lol Check all that apply. 12x2 - 2x - 2 = 2 A. x = 7 and x = 3 B. x = 7/3 and x = -3/2 C. x = 2/3 and x = -1/2 D. x = 7/3 and x = 3/2

sam (.sam.):

3x2 + 7x + 8 = 6 3x2 + 7x + 2 = 0 Factor (x+2)(3x+1)=0 x+2=0 x=-2 ---------- or ---------- 3x+1=0 x=-1/3 x=-2 , -1/3

OpenStudy (anonymous):

x=-1/2,2/3

OpenStudy (anonymous):

THANK YOU SAM :)

sam (.sam.):

welcome :)